OpenStudy (anonymous):

substitute 4 for x and find the matching value for y. x-6y=-2

OpenStudy (ahsome):

Ok, your equation is \[x-6y=-2\]Right?. Now, it says that \(x\) is 4. So, whenever there is an \(x\), put a 4 instead \[x-6y=-2\]\[4-6y=-2\]

OpenStudy (ahsome):

Now, we only want the 4 to be by itself. So, since the 4 is positive, we need to add a negative 4 on both sides \[4-6y=-2\]\[4-6y-4=-2-4\] The reason we add on both sides is so the equation is balance. If we keep it only on one, it would change the value We can cancel the two fours on the left hand side. Because \(4-4=0\) \[4-6y-4=-2-4\]\[\cancel{4}-6y\cancel{-4}=-2-4\]\[-6y=-2-4\] \(-2-4=-6\), Right? \[-6y=-2-4\]\[-6y=-6\]. Now, divide both sides by \(-6\), to get rid of it on the left hand side \[-6y=-6\] \[\frac{-6y}{-6}=\frac{-6}{-6}\]\[\frac{\cancel{-6}y}{\cancel{-6}}=\frac{\cancel{-6}}{\cancel{-6}}\] \[y=1\] Tada :)

OpenStudy (anonymous):

thank you!

OpenStudy (ahsome):

No problem :) @srodda15